OpenStudy (anonymous):

What is n/5 - 3n/10 = 1/5

OpenStudy (anonymous):

answer is -2

OpenStudy (anonymous):

n/5 -3n/10 taking lcm of 5 and 10 you get 10 so now \[\frac{ 2n-3n}{ 10 }\] -n/10 -n/10=1/5 -n=2 n= -2 hope you get it @Fifiluv101